OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 5 hours, helped by a wind of 2 miles per hour. Pedaling at the same rate, the trip back against the wind takes 9 hours. Find find the total round trip distance she traveled.

OpenStudy (anonymous):

Let x be the speed of the bike in no wind. r*t=d Solve the following for x: (x + 2) 5 = 9 (x - 2) x=7 mph Using the LHS of (x + 2) 5 = 9 (x - 2) evaluated at x=7. (7 + 2) 5=45miles The round trip distance is 2*45 = 90 miles.